OpenStudy (anonymous):

y = 2x - 4 a. (6,8) b. (1, -4) c. (-2, -5)

OpenStudy (anonymous):

can A be correct?you say first...

OpenStudy (anonymous):

Look : we can change that and write this : y-2x=-4

OpenStudy (anonymous):

8 = (2x6) - 4 is this possible?

OpenStudy (anonymous):

yes! 8 = 12 -4 = 8

OpenStudy (anonymous):

so A is correct,

OpenStudy (anonymous):

-4 = (2x1) -4

OpenStudy (anonymous):

it's not correct...

OpenStudy (anonymous):

-5 = (-2x2) -4

OpenStudy (anonymous):

it's not correct too,so A is correct, got it?

OpenStudy (anonymous):

@PFEH.1999 Yes I did thanks

OpenStudy (anonymous):

your welcome ;)